# Self-Reflection in LLM Agents: Effects on Problem-Solving Performance

### v2

@@ 1.2 Prior Literature
However, these LLM agents have several imitations. They have limited knowledge, make errors in reasoning, hallucinate output, and get stuck in unproductive loops [4, 5, 6, 7, 8, 9].
@@
Self-reflection in LLM agents is a meta-cognitive strategy also known as introspection [13, 14]. Some research studies have indicated that LLMs using self-reflection are able to identify and correct their mistakes [12, 10, 8, 15]. Others have indicated that LLMs cannot identify errors in their reasoning; regardless, they still may be able to correct them with external feedback [7, 26].
@@ 1.3 Contribution
In addition, it provides insight into which types of LLMs and problem domains benefit most from each type of self-reflection. These include LLMs like GPT-4, Llama 2 70B, and Gemini 1.5 Pro. It also includes various problem domains such as math, science, medicine, etc.
@@ 2.1 Data
We preprocessed and converted these datasets into a standardized format. Then, we randomly selected 100 questions from each of the ten datasets to create a multi-domain exam with 1,000 problems.
@@ 2.2 Models
We evaluated our agents using nine popular LLMs, including GPT-4, Llama 2 70B, Google Gemini, etc. [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. All models were accessed via cloud-based APIs hosted by Microsoft, Anthropic, and Google.
@@
Mistral Large Mistral AI 2024-02-26 Closed [47]
@@ 2.4 Process
Then, a find-and-replace operation was performed on the text of each self-reflection to redact the answer labels and answer descriptions. For example, we replaced answer labels (e.g., "A", "B", "C") and answer descriptions (e.g., "Baltimore", "Des Moines", "Las Vegas") with the text "[REDACTED]". [^1] The process we used to redact answer labels and descriptions was greedy. It often redacting additional text that did not leak the answer. However, we felt it necessary to err on the side of caution by eliminating any possible answer leakage. This was done to all of the self-reflecting agents, except for the Unredacted agent, to prevent answer leakage in the self-reflections. [^2] It is important to note that the self-reflections generated by the Explanation, Instructions, and Solution agents indirectly leak information about the correct answer without directly specifying the correct or incorrect answers. However, they generated this information on their own based on nothing more than being provided the correct answer during the self-reflection process.
@@ 2.5 Metrics
We used correct-answer accuracy as our primary metric to measure the performance of the agents. We divided the number of correctly answered questions by the total number of questions.
@@ 3.2 Performance by Model
In terms of performance by model, every LLM that we tested demonstrated similar increases in accuracy across all self-reflection types. In all cases, the improvement in performance was statistically significant ($p<0.001$). See Figure 3 for a plot of accuracy by model and agent. See Table 4 for a numerical analysis of accuracy across all models.
@@ 4.1 Interpretation
The difference in accuracy between the self-reflecting agents and the Unredacted agent demonstrated that we were effectively eliminating direct answer leakage from the self-reflections. However, the structure of feedback generated by the Instruction, Explanation, Solution, and Composite agents clearly provides indirect guidance toward the correct answer without directly giving the answer away.
@@ 4.2 Limitations
Second, API response errors may have introduced a small amount of error into our results. API errors typically occurred when content-safety filters were triggered by the questions being asked. In most cases, this may have amounted to an error in reporting an agent’s accuracy of less than 1%. However, in the case of the Gemini 1.0 Pro and the Mistral Large models, this error could be as high as 2.8%.
Third, the top-performing LLMs scored above 90% accuracy for most exams. As a result, the increase in scores for the top exams was compressed near the upper limit of 100% (i.e., a perfect score). This compression effect makes it difficult to assess the performance increase. As a result, our analysis would benefit from exams with a higher level of difficulty.
@@ 4.3 Implications
Our research builds upon previous research in LLM agents and self-reflecting agents.
@@
In addition, our research has theoretical implications for AI researchers studying meta-cognition in LLMs. If LLMs are able to self-reflect on their own CoT, other similar metacognitive processes may also be leveraged to improve their performance.
@@ 5 Conclusion
In this study, we investigated the effects of self-reflection in LLM agents on problem-solving tasks. Our results indicate that LLMs are able to reflect upon their own CoT and produce guidance that can significantly improve problem-solving performance. These performance improvements were observed across multiple LLMs, self-reflection types, and problem domains. This research has practical implications for AI engineers building agentic AI systems as well as theoretical implications for AI researchers studying meta-cognition in LLMs.
@@ 6 Acknowledgements
Funding for this research was provided by Microsoft and Renze AI Research Institute.
@@ A.1 Results
Agent Title AQUA-RAT ARC Hellaswag LSAT-AR LSAT-LR LSAT-RC LogiQA MedMCQA SAT-English SAT-Math

### v3

@@ 1.2 Prior Literature
However, these LLM agents have several limitations. They have limited knowledge, make errors in reasoning, hallucinate output, and get stuck in unproductive loops [4, 5, 6, 7, 8, 9].
@@
Self-reflection in LLM agents is a metacognitive strategy also known as introspection [13, 14]. Some research studies have indicated that LLMs using self-reflection are able to identify and correct their mistakes [12, 10, 8, 15]. Others have indicated that LLMs cannot identify errors in their reasoning; regardless, they still may be able to correct them with external feedback [7, 26].
@@ 1.3 Contribution
In addition, it provides insight into which types of LLMs and problem domains benefit most from each type of self-reflection. These include LLMs such as GPT-4, Llama 2 70B, and Gemini 1.5 Pro. It also includes various problem domains such as math, science, medicine, etc.
@@ 2.1 Data
We preprocessed and converted these datasets into a standardized format. Then, we randomly selected 100 questions from each of the ten datasets to create a multi-domain exam consisting of 1,000 problems.
@@ 2.2 Models
We evaluated our agents using nine popular LLMs, including GPT-4, Llama 2 70B, Gemini 1.5 Pro, etc. [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. All models were accessed via cloud-based APIs hosted by Microsoft, Anthropic, and Google.
@@
Mistral Large Mistral AI 2024-02-26 Open [47]
@@ 2.4 Process
Then, a find-and-replace operation was performed on the text of each self-reflection to redact the answer labels and answer descriptions. For example, we replaced answer labels (e.g., "A", "B", "C") and answer descriptions (e.g., "Baltimore", "Des Moines", "Las Vegas") with the text "[REDACTED]". [^1] The process we used to redact answer labels and descriptions was greedy. It often redacted additional text that did not leak the answer. However, we felt it necessary to err on the side of caution by eliminating any possible answer leakage. This was done to all of the self-reflecting agents, except for the Unredacted agent, to prevent answer leakage in the self-reflections. [^2] It is important to note that the self-reflections generated by the Explanation, Instructions, and Solution agents indirectly leak information about the correct answer without directly specifying the correct or incorrect answers. However, they generated this information on their own based on nothing more than being provided the correct answer during the self-reflection process.
@@ 2.5 Metrics
We used correct-answer accuracy as our primary metric to measure the performance of the agents. Accuracy is calculated by dividing the number of correctly answered questions by the total number of questions.
@@ 3.2 Performance by Model
In terms of performance by model, every LLM that we tested demonstrated similar increases in accuracy across all types of self-reflection. In all cases, the improvement in performance was statistically significant ($p<0.001$). See Figure 3 for a plot of accuracy by model and agent. See Table 4 for a numerical analysis of accuracy across all models.
@@ 4.1 Interpretation
The difference in accuracy between the self-reflecting agents and the Unredacted agent demonstrates that we were effectively eliminating direct answer leakage from the self-reflections. However, the structure of feedback generated by the Instruction, Explanation, Solution, and Composite agents clearly provides indirect guidance toward the correct answer without directly giving the answer away.
@@ 4.2 Limitations
Second, API response errors may have introduced a small amount of error into our results. API errors typically occurred when content-safety filters were triggered by the questions being asked. In most cases, this may have amounted to an error in reporting an agent’s accuracy of less than 1%. However, in the case of the Gemini 1.0 Pro and Mistral Large models, this error could be as high as 2.8%.
Third, the top-performing LLMs scored above 90% accuracy for most exams. As a result, the increase in scores for the top exams was compressed near the upper limit of 100% (i.e., a perfect score). This compression effect makes it difficult to accurately assess the performance increase. As a result, our analysis would benefit from exams with a higher level of difficulty.
@@ 4.3 Implications
Our research builds upon prior work on LLM agents and self-reflection.
@@
In addition, our research has theoretical implications for AI researchers studying metacognition in LLMs. If LLMs are able to self-reflect on their own CoT, other similar metacognitive processes may also be leveraged to improve their performance.
@@ 5 Conclusion
In this study, we investigated the effects of self-reflection in LLM agents on problem-solving tasks. Our results indicate that LLMs are able to reflect upon their own CoT and produce guidance that can significantly improve problem-solving performance. These performance improvements were observed across multiple LLMs, self-reflection types, and problem domains. This research has practical implications for AI engineers building agentic AI systems as well as theoretical implications for AI researchers studying metacognition in LLMs.
@@ 6 Acknowledgements
Funding for this research was provided by Microsoft and the Renze AI Research Institute.
@@ A.1 Results
Agent Name AQUA-RAT ARC Hellaswag LSAT-AR LSAT-LR LSAT-RC LogiQA MedMCQA SAT-English SAT-Math